OpenStudy (anonymous):

Is -18 great than or equal to 5 x - 8

OpenStudy (texaschic101):

Is your question supposed to be : -18 <= 5x - 8, and your supposed to find x ?

OpenStudy (texaschic101):

Because if you supposed to find out if -18 is <= to 5x - 8, then we need to know the value of x to determine if -18 is greater then or= to.

OpenStudy (texaschic101):

are you supposed to solve for x in the inequality ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

-18 <= 5x - 8 (add 8 to both sides) -18 + 8 <= 5x (combine like terms) -10 <= 5x (now divide both sides by 5) -2 <= x

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No. Thank you

OpenStudy (texaschic101):

anytime :)